A production line produces 15 CD players every 4 hr. How many players can it produce in 39 hr?
(A) What kind of variation applies to this situation?
(B) Solve the problem.
What kind of variation applies to this situation?
1. Direct variation
2. Inverse variation
In 39 hr, the production line can produce how many CD players?
___ CD players

Answer by stanbon(75887)

A production line produces 15 CD players every 4 hr. How many players can it produce in 39 hr?
(A) What kind of variation applies to this situation?:: direct
p = k*h
15 = k*4
k = 15/4
----
Equation:
p = (15/4)h
Ans: p(39) = (15/4)39 = 150 players when rounded up
-------------------------
Cheers,
Stan H.
-----------